OpenStudy (anonymous):

What is the probability that a state will get hit by a major tornando at least once in the next 10 years if the probability of getting hit in one year is 1/10?

OpenStudy (anonymous):

poisson distribution for this one \[\lambda=0.1\]

OpenStudy (anonymous):

and "at least once" means "not none" so caclulate the probability of no tornadoes and subtract the result from one

OpenStudy (anonymous):

actually i think it is easier than that. since you are using a ten year period and the probability is 1/10 for one year, you can use \[\lambda =1\]

OpenStudy (anonymous):

\[P(x=0)=e^{-\lambda}=e^{-1}\]

OpenStudy (anonymous):

does this look familiar or are you supposed to do it a different way?

OpenStudy (anonymous):

i think your answer is \[1-e^{-1}\]

OpenStudy (anonymous):

I think a different way but I'm not sure...sorry

OpenStudy (anonymous):

are you using the poisson distribution? i am not sure how else you could do this one

OpenStudy (anonymous):

i guess you could say they are independent, so probability of no tornadoes in any one years is \[\frac{9}{10}\] so no tornadoes 10 years in a row is \[(\frac{9}{10})^{10}\] so "at least one" in ten years is \[1-(\frac{9}{10})^{10}\] that might work better

OpenStudy (anonymous):

using that method i get 0.65 as your probability

OpenStudy (anonymous):

yeah that definitely looks more like what im doing haha

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

poisson gives 0.63 but that is more of an approximation, so stick with that one